OpenStudy (anonymous):

given a leg with 14 km and a hypotenuse 19 km, find the missing side

OpenStudy (anonymous):

use pythagorean thereom: a^2 + b^2 = c^2 so: 14^2 + b^2= 19^2 simplify: 196 + b^2 = 361 isolate b^2: b^2=165 isoalte b: 12.84523257866513